• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name real j.s.

    Log entry time 16:09:09 on June10,2002

    Entry number 85772

    keyword=shift summary

    Kinematics Q1_l40 start at run 1368.
    8:06 we start the run 1371
    8:13 acquisition dead time is 20%. we will change the prescale.
    run 1372 dead time is still 20% with a rate of 1200 evts/sec
    should be ~800 from reading xscaler!!
    9:15 MCC call. No beam for a while... stop the run 1374.
    9:47 beam back. try an harp scan: does not work
    9:55 go back data taking run 1375
    all rates before this run have been read on helicity - from xscaler
    this explain the difference in counting rates we had between coda
    and xscaler. prescale have been adjust.
    10:54 mcc will reboot iocha1. run 1377 has been stopped
    11:05 mcc ask us to reboot the ioc of the target
    11:15 harp scan run is 1378
    11:25 go back production run 1379
    12:15 the readback for the helium flow for Q1 left shows -23.5. we call scot but
    since there is no trip let's continue to take data. It seems to be readout
    problem. Scot et al. will investigate this problem and may need an entry.
    this will be done during the next maintenance day
    15:10 mcc call, they will change the fan of vme crate. beam is off for a while
    15:45 beam back
    16h00 last run is 1388 kinamatics Q1_l40. still 1hour of running this